For the function y =x-a/a^2-b^2,
and the initial value of the independent variable x0 = a - b. At what terminal value of x the
increment ?y =1/a-b

Grade:upto college level

1 Answers

Sher Mohammad IIT Delhi
10 years ago
y=x-a/(a+b)(a-b)
a-b=x-a/y(a+b)
at x=-b , y=-(a+b)/(a+b)(a-b), y=-1/(a-b)

sher mohammad
b.tech, iit delhi